OpenStudy (anonymous):

If the clothing maker bought 500 m2 of this fabric, how much money did he lose? Use 1tepiz=0.625dollar and 0.9144m=1yard.

OpenStudy (theeric):

Hi! Is that the whole question?

OpenStudy (anonymous):

this is the top im not sure if this matters: Here is a simple example. The inhabitants of a small island begin exporting beautiful cloth made from a rare plant that grows only on their island. Seeing how popular the small quantity that they export has been, they steadily raise their prices. A clothing maker from New York, thinking that he can save money by "cutting out the middleman," decides to travel to the small island and buy the cloth himself. Ignorant of the local custom of offering strangers outrageous prices and then negotiating down, the clothing maker accepts (much to everyone's surpise) the initial price of 400 tepizes/m2. The price of this cloth in New York is 120 dollars/yard2.

OpenStudy (theeric):

Ah, I see! Yeah. How this works is, this guy paid 400 tepiz/m^2 and recieved $12/yd^2. Overall, he lost 400 tepiz/m^2 times the square meters, and gained $12/yd^2 times the square yards of fabric. Are you okay with that much? Let me know if you have any questions!

OpenStudy (theeric):

Well, \(400\text{ tepiz}/m^2 * 500\ m^2 = 20,000\text{ tepiz}\) That is what he lost. Now you need to convert if you want it in dollars!

OpenStudy (theeric):

You want a conversion factor, which is just - you multiply something by a conversion factor to change the units without changing the amount! Same amount, different units. So, since it doesn't change the amount when you multiply it, the conversion factor must be equal to \(1\). How? You are given that 1tepiz=0.625dollar. Divide both side by \(1\text{ tepiz}\) to get \(\dfrac{1\text{ tepiz}}{1\text{ tepiz}}=1=\dfrac{$0.625}{1\text{ tepiz}}\). So you have that. Now, look at how it helps. \(200,000\cancel{\text{ tepiz}}\times \dfrac{$0.625}{1\cancel{\text{ tepiz}}}=$125,000\) He spent $125,000. Now how much did he make? You'll have to do something similar to find out how much! :)